OpenStudy (anonymous):

use de moivre's theorem to write the complex number in trigonometric form. ((sqrt3)+i)^3

OpenStudy (kropot72):

Do you know how to put the complex number into polar form:\[r(\cos \theta+i \sin \theta)\]

OpenStudy (anonymous):

ya

OpenStudy (kropot72):

So what is the polar form of the complex number \[\sqrt{3}+i\]